Suppose the market demand and market supply curves are given by the equations:

Qd= 100-P

Qs= 3P

a. What are the equilibrium price and equilibrium quantity in the market for this product?

b. Find out consumer surplus, producer surplus, and total surplus.

c. Suppose the government sets a price floor at $26 for this product. With this price floor, how much is consumer surplus?

d. With this price floor of $26, how much is producer surplus?

e. Find out total surplus and deadweight loss(DWL).
Answer #1

A) Set D=S

100-P = 3P

100 = 3P+P

100 = 4P

P = 100/4 = 25

Q = 100-25 = 75

B) Maximum reservation price = 100

Minimum reservation price = 0

CS = 0.5*75*(100-25) = 2812.5

PS = 0.5*75*(25-0) = 937.5

TS = CS+PS = 3750

C) With PF = 26

QD = 100-26 = 74

CS = 0.5*74*(100-26) = 2738